OpenStudy (anonymous):

I need help please. What is 2xy*3x^2y? Thankyou very much mateys.

OpenStudy (lgbasallote):

do you know the value of these: 2 * 3 x * x^2 y * y ??

OpenStudy (anonymous):

6, ermmmm, y^2. I dont know. It has been a long holiday.

OpenStudy (lgbasallote):

right...what about x * x^2?

OpenStudy (anonymous):

x^3?

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

Aw thanks!

OpenStudy (lgbasallote):

2*3 = 6 x * x^2 = x^3 y* y = y^2 so put all of those together...

OpenStudy (lgbasallote):

what do you get?

OpenStudy (anonymous):

6+x^3+y^2

OpenStudy (lgbasallote):

nope..not plus

OpenStudy (anonymous):

Arrrr

OpenStudy (lgbasallote):

see how in the original equation they are just "stuck to one another"

OpenStudy (anonymous):

times?

OpenStudy (lgbasallote):

yep

OpenStudy (anonymous):

Ok thanks!

OpenStudy (lgbasallote):

welcome

OpenStudy (lgbasallote):

sure